OpenStudy (anonymous):

What is the slope of a line parallel to the line with equation 5x – 2y = 11?

OpenStudy (♪chibiterasu):

Well since both lines will have the same slope, all you have to do is convert his into slope intercept and the slop of this line will be the slope of the line parallel to it. Do you know what slop intercept form looks like?

OpenStudy (solomonzelman):

Subtract 5x from both sides and divide both sides by '-2'. What is the slope? Parallel lines always have the same slope.

OpenStudy (♪chibiterasu):

Well, slope intercept form is: y = mx + b m is the slope so whatever is plugged into m will be your answer.

OpenStudy (♪chibiterasu):

And yeah, thanks >_>

OpenStudy (anonymous):

A. -11/2 B. -5/2 C. -2/5 D. 5/2

OpenStudy (♪chibiterasu):

Well, lets subtract 5x from both sides: 5x - 5x – 2y = 11 - 5x -2y = 11 - 5x Now we just divide -2 on both sides: -2y/-2 = -11/2 - 5x y = -11/2 - 5x We know that m is the slope so -11/2 is our answer.

OpenStudy (anonymous):

Which of the following has a slope of –3 and a y-intercept of –4?

OpenStudy (anonymous):

y = –3x + 4 y = 3x + 4 y = 3x – 4 y = –3x – 4

OpenStudy (♪chibiterasu):

Well, y = mx + b m is the slope and b is the y intercept.

OpenStudy (♪chibiterasu):

So I think you can do this on your own knowing m and b :3

OpenStudy (anonymous):

b is the asnwer?

OpenStudy (♪chibiterasu):

Um, noooo . ..

OpenStudy (♪chibiterasu):

Oh thats what you emant

OpenStudy (solomonzelman):

y = 3x + 4 compare to y=mx+b

OpenStudy (anonymous):

cause i thought it was C?

OpenStudy (♪chibiterasu):

No, try again

OpenStudy (♪chibiterasu):

Both the slope and y-intercept are negative o:

OpenStudy (solomonzelman):

m is the slope b is the y-intercept y=mx+b

OpenStudy (♪chibiterasu):

Yes, I mentioned that like 5 times already.

OpenStudy (solomonzelman):

–3 is the slope. -4 is the y-intercept.

OpenStudy (anonymous):

oh im sorry was looking at the wronmg one its D

OpenStudy (solomonzelman):

Yup!

OpenStudy (♪chibiterasu):

Hey, thanks blurting the answer when I wanted her to do it on her own ;-;

OpenStudy (♪chibiterasu):

But yes its D :c

OpenStudy (solomonzelman):

Did I say the answer?

OpenStudy (anonymous):

Which ordered pair is on the graph of the equation 4x + 3y = 13?

OpenStudy (anonymous):

(1, 6) (1, –3) (0, –3) (4, –1)

OpenStudy (solomonzelman):

Put the '4x + 3y = 13' into slope intercept form. Can you do that?

OpenStudy (solomonzelman):

Actually just plug in points to see which fits and which doesn't.